OpenStudy (anonymous):

someone help please What is the x-intercept of the line with the equation? 1/3x + y = -27

OpenStudy (korosh23):

Here x is a unit I believe. Am I right?

OpenStudy (anonymous):

Solve the equation for y and the set x to 0

OpenStudy (anonymous):

yea @korosh23 & ok @TheDream

OpenStudy (anonymous):

What answer do you get?

OpenStudy (korosh23):

Well this has a way ti solve.

OpenStudy (korosh23):

1/3x + y= -27 x + y= - 27(3) x= -81- y and y= -81 - x